The mass contributions of the two isotopes are:

61.27 u Rb-85

24.20 u Rb-87

What is the average atomic weight?
1
Expert's answer
2020-05-04T01:52:25-0400

average atomic weight = "\sum" atomic weight of the isotopes * mass contributions of the isotopes


According to the conditions of the problem, average atomic weight of Rb = 85*0.6127 + 87*0.2420 = 73.13 u.


But these values of mass contributions are wrong! The correct values (taken from Wiki) are 72.2% for Rb-85 and 27.8% for Rb-87. Use these values for calculations:


Average atomic weight of Rb = 85*0.722 + 87*0.278 = 85.56 u.